### v2

# Evidence of mini-jet emission in a large emission zone from a magnetically-dominated gamma-ray burst jet

Abstract
The second brightest GRB in history, GRB230307A provides an ideal laboratory to study the details of GRB prompt emission thanks to its extraordinarily high photon statistics and its single broad pulse overall shape characterized by an energy-dependent fast-rise-exponential-decay (FRED) profile. Here we demonstrate that its broad pulse is composed of many rapidly variable short pulses, rather than being the superposition of many short pulses on top of a slow component. Such a feature is consistent with the picture of many mini-jets due to local magnetic reconnection events in a large emission zone far from the GRB central engine, as envisaged in the internal-collision-induced magnetic reconnection and turbulence (ICMART) model, but raises a great challenge to the internal shock models that attribute all variability components to collisions among different shells. Since relativistic mini-jets demand strong magnetization in the outflow, this work provides strong evidence for a Poynting-flux-dominated jet composition of this bright GRB.
Key Laboratory of Particle Astrophysics, Institute of High Energy Physics, Chinese Academy of Sciences, Beijing 100049, China.
University of Chinese Academy of Sciences, Chinese Academy of Sciences, Beijing 100049, China.
Department of Astronomy, Beijing Normal University, Beijing 100875, People’s Republic of China.
Nevada Center for Astrophysics, University of Nevada Las Vegas, NV 89154, USA.
Department of Physics and Astronomy, University of Nevada Las Vegas, NV 89154, USA.
ICRANet, Piazza della Repubblica 10, I-65122 Pescara, Italy
ICRA, Dipartamento di Fisica, Sapienza Università di Roma, Piazzale Aldo Moro 5, I-00185 Rome, Italy
INAF, Osservatorio Astronomico d’Abruzzo, Via M. Maggini snc, I-64100, Teramo, Italy
College of Electronic and Information Engineering, Tongji University, Shanghai 201804, China
College of Physics and Hebei Key Laboratory of Photophysics Research and Application, Hebei Normal University, Shijiazhuang, Hebei 050024, China
Guizhou Provincial Key Laboratory of Radio Astronomy and Data Processing, Guizhou Normal University, Guiyang 550001, China
School of Physics and Optoelectronics, Xiangtan University, Yuhu District, Xiangtan, Hunan, 411105, China
School of Computer and Information, Dezhou University, Dezhou 253023, China
The origin of the prompt emission of gamma-ray bursts (GRBs) is still subject to debate because of the not-well-constrained jet composition, location of the emission region, and mechanism with which -rays are produced[1]. Depending on the unknown jet composition, three emission sites are commonly discussed in GRB prompt emission models. For a matter-dominated fireball, the observed emission is likely a superposition of a thermal component originating from the fireball photosphere at cm[2, 3, 4] and a non-thermal component orignating from synchrotron-radiating electrons accelerated from internal shocks at cm[5, 6, 7]. Within these models, rapid variability in GRB prompt emission lightcurves is attributed to erratic central engine activities, and the emission radius is no larger than the internal shock radius defined by , where is the variability timescale of the short pulses, and is the average bulk Lorentz factor of the jet. The non-detection of bright photosphere emission[8, 9] and non-detection of high-energy neutrinos from any GRBs[10, 11] strongly suggested that a large fraction of GRBs do not have a matter-dominated composition. Rather, the outflows are likely Poynting-flux-dominated even at radii way above the photosphere. Within such a picture, the emission from the photopshere is suppressed. If the magnetization factor is in the emission region, mini-jets may be emitted[12, 13, 14] because magnetic reconnection events would drive a local relativistic flow with Lorentz factor . A well-developed model in this regime is the internal-collision-induced magnetic reconnection and turbulence (ICMART) model[14], which envisages an emission radius cm, where is the timescale of the broad pulses in GRB lightcurves. Within this model[15, 16], short pulses with timescale do not correspond to individual shell collisions, but are rather produced by individual reconnection events occuring at a large emission site defined by . The emission in the timescale of is expected to follow the single-shell evolution and displays an energy-dependent behavior such as energy-dependent duration, spectral lags and global hard-to-soft evolution[17, 18]. The emission in the timescale of , on the other hand, reflects the behavior of individual reconnection events, and is not expected to display all those features.
The extremely bright GRB 230307A was firstly reported by the Gravitational wave high-energy electromagnetic counterpart all-sky monitor (GECAM)[19] with trigger time of 15:44:06.650 UT on 7 March 2023 (denoted as T). The Lobster Eye Imager for Astronomy (LEIA, the pathfinder of the Einstein Probe mission) also observed its prompt emission in the soft X-ray band[20]. While the burst’s duration (measured as ) of about 41 seconds aligns with the category of long-duration GRBs, the association of a kilonova[21] and its unique properties in the prompt emission strongly suggest its origin from a compact binary merger event[20].
Despite of the extraordinary brightness, unsaturated data record by GECAM enable us to accurately characterize the temporal properties of GRB 230307A (see Methods). The GECAM time-binned net count rate as a function of time in different energy bands (multi-band light curves hereafter) show a typical FRED profile (Figure 1a). Therefore, we use a conventional parameterization of the FRED profile [22] (“Norris05" hereafter)
to fit the light curves in each energy bands. In the above formulation, is the starting instance of the pulse, and are the rising and decaying time scales, respectively. The fitting results are listed in Table 1. The peak time and the width of the pulse are therefore defined as: and . It can be seen in Figure 2 that there is a clear energy dependence in both and , as commonly observed in GRB broad pulses. In the energy range from 6 keV to 300 keV, the and relations are both in a power law function, with the power index -0.38 and -0.37, respectively. Above keV, the energy dependence on both and saturate. This highly correlated and relationship implies self-similar profiles across different energy bands, a feature that has also been reported for other bursts in the literature[22, 23, 24, 25]. For the details of the broad feature fitting see Methods.
[…]
a, The black histograms are 1 s binned net count rates (with background subtracted), whose profile shows a single FRED shape; The gray thin histograms are 0.1 s binned net count rates (with background subtracted); The blue solid curves are best-fit FRED model with the Norris05 formulation, while the blue dashed curves are the best-fit with our new FRED formulation. Gray shadowed regions (the precursor and the dip) are ignored in the fitting. All error bars represent 1 uncertainties of the net count rates.
b, The residuals of the net light curves after fitting with the new FRED formulation.
c, The cross-correlation between the multi-band residuals and that in the 6-30 keV channel. The vertical dotted line indicates the zero lag time.
In order to further demonstrate the self-similar feature of the slow varying profile, we smooth the light curves in a non-parametrical way, and manifest their shape identity after energy-dependent re-scaling in time (see Methods). We stack all the re-scaled and smoothed light curves in the Extended Data Fig 1a. It can be obviously shown that, the shape of the profiles in different energy bands can be "stretched" in time domain into an identical shape with an energy dependent scaling factor. The scaling factor as a function of energy is plotted in the Extended Data Fig 1b.
The self-similarity inspired us to propose another formulation of FRED profile, with one less parameter than that of Norris05:
We fit the multi-band light curves again with the new FRED formulation. The results of fitting are listed in Methods. The sole time scale parameter shows a similar energy dependence on energy, , from 6 to 300 keV, and shows a hint of shallower energy dependence above keV (Figure 2). In the new formulation, one can define peak time as , where we denote the in the channel with . The width of the FRED profile also scales as . Therefore, the found dependence can naturally result in the and relations.
Beside the self-similar FRED profile of the broad feature of GRB0307A’s light curve, the extraordinary brightness of the burst allows us to study the fast varying temporal structures. From the residuals of the light curves (Figure 1b), one can see that there are many rapidly-varying short pulses. One noticeable feature is that even the broad pulse has a clear energy dependence, the short time spikes and dips appear to align at the same instances across different energy bands. We further demonstrate this alignment of the fast temporal features by cross-correlations among multi-band light curve residuals. As shown in the Figure 1c, the peaks are perfectly aligned across the full energy band. This immediately excludes the attempt to explain the self-similar stretchable feature of the multi-band light curves as relativistic time-dilation, as a time-dilation effect would stretch both slow and fast varying pulses together.
We show that some of these prominent short duration structures can be fitted with individual short pulses without an underlying slow component(see Methods). The negative residuals, most prominently the dip at around 18 s in all energy bands, essentially excludes the possibility that short time structures are added onto a broad pulse component. It is also difficult to explain this dip as absorption, since the effective optical depth shows an energy dependence which no known absorption mechanism can reproduce (see Methods). Beside, if we were to explain the dip as absorption, its depth and transience require that the absorbing object has a angular size comparable to that of the emission region, and a very high transverse angular velocity. That will either result in a superluminal transverse velocity of the blocking object, or highly coincidental absorption in our local. The most likely explanation is that, the broad pulse of the light curves is composed of many fast pulses, and the dips are gaps between successive short pulses (see Methods).
The aforementioned results that (a) the overall light curve resembles a classic single pulse, which is wider and peaks later in lower energies in a self-similar way; and that (b) there is no underlying elementary single slow pulse and the apparent broad pulse is composed with many fast pulses***We note that even though the feature (a) has been observed and carefully studied before in several other bursts (e.g. the first pulse of GRB130427A[26]), the feature (b) is ambiguously unveiled for the first time only in this burst., serve as important tests on the GRB prompt emission models. The standard internal shock model attributes short-time pulses as a consequence of collisions of many pairs of shells (see Extended Data Fig. 4 for an illustration). The typical emission region is at[5, 6] from the central engine, where is the bulk Lorentz factor of the GRB, and s is the typical variability time scale of the fast component. Within such a scenario, the broad pulse is defined as the history of the central engine activity. Lightcurves in different bands should share this same history and should not display a global energy-dependent behavior (see Methods: Physical implications for detailed simulations that justify this claim). The observed clear energy-dependence of the broad pulse abandons this interpretation. A modified internal shock model (see Extended Data Fig. 5) may invoke another internal shock at a much larger emission radius defined by , where s is the timescale of the broad pulse. However, within such a picture, the observed emission should be the superposition between the fast and slow components, which requires that fast pulses only grow on top of a background flux defined by the slow component. The very deep, achromatic dip at around 18 s, therefore, rules out such a possibility. We therefore conclude that the internal shock models are clearly disfavored. Since all types of photosphere models invoke emission radius smaller or equal to , they are also disfavored by the same reasoning.
The ICMART model, on the other hand, offers a natural interpretation to the data. Within this model[14, 15, 16], the slow component with a characteristic timescale is related to the typical thickness of the magnetic blobs. The collision site between two magnetically-dominated shells is at . The rapid-varying fast component with timescale , on the other hand, is related to the emission timescale of local mini-jets within the global emission zone (see illustration in Extended Data Fig. 6). The broad pulse is simply the superposition of emission of many mini-jets. No underlying slow component is needed. Significant dips are allowed, especially during the decay phase when high-latitude emission starts to play a role. The location of the dip is consistent with such a phase[20]. Because the fast pulses originate from local mini-jet events, the broad-band emission of the fast component is related to the dynamics of the mini-jets and therefore perfectly aligned in time, as is observed.
The general trend of the energy-dependent peak time and duration is naturally expected within the framework of the ICMART model. Unlike the internal shock model that attributes different short pulses as emission from different emitting fluids, the ICMART model suggests that the emission of the entire broad pulse originates from the same fluid as it expands in space. The magnetic field strength in the emission region decays with radius due to expansion[27], leading to rolling down of the characteristic synchrotron emission frequency with time. As result, higher energy emission peaks earlier and lasts shorter than lower energy emission[17, 18]. The general trend of model prediction matches the observations well[17] (see Methods: Physical implications for detailed simulations within the framework of the ICMART model that justify our claim), even though the self-similarity revealed from the data may require special model parameters, which may be related to the details of magnetic turbulence and reconnection in the emission region, and thus provide new insights into the physics of gamma-ray bursts. It is interesting to note that within this picture, the observed prompt emission originates from a single episode of central engine activity. The long-duration of the burst is related to the spreading of the magnetic blob at large radii. This effect reduces the intrinsic duration of the central engine activity, which makes the burst more consistent with the merger origin of GRB 230307A[28, 20, 29].
1 Method
1.1 Observations
GECAM is a dedicated all-sky gamma-ray monitor constellation funded by the Chinese Academy of Sciences, and now consists of three telescopes, i.e. GECAM-A and GECAM-B[30] micro-satellties launched together on December 10, 2020, and GECAM-C (also called High Energy Burst Searcher, HEBS)[31] onboard SATech-01 experimental satellite launched on July 27, 2022. There are two kinds of detectors in each GECAM telescope: Gamma-Ray Detectors (GRDs) and Charged Particle Detectors (CPDs). GRDs are the main detector of GECAM, each of which is composed of a scintillator and an array of SiPMs. There are 25 GRDs onboard each of GECAM-A and GECAM-B, and 12 GRDs onboard GECAM-C. GECAM is designed to have the highest time resolution (0.1 s) among all GRB detectors ever flown[32], making it one of the most suitable detectors for conducting the temporal research about GRBs.
GRB 230307A triggered GECAM-B in real-time at 15:44:06.650 UT on 7 March 2023 (denoted as T) and also detected by GECAM-C and other gamma-ray monitors (e.g. Fermi/GBM, Konus-Wind), while GECAM-A was offline at that time. The real-time alert data was transmitted instantly with the Global Short Message Communication of Beidou satellite navigation system[33, 34], and processed by the automatic pipeline of GECAM[35], based on which the extreme brightness of GRB 230307A was firstly reported by GECAM to the community[19], initiating many multi-wavelength follow-up observations[36, 37]. GRB 230307A is specially interesting because of its apparent single FRED shape in the overall light curve, and its high fluence and variability.
Thanks to the dedicated design of instrument[31, 38], neither GECAM-B nor GECAM-C suffered from data saturation during the whole burst of GRB 230307A despite of its extreme brightness[20]. High quality of GECAM data allow us to accurately measure the temporal and spectral properties of GRB 230307A. GRD04 of GECAM-B and GRD01 of GECAM-C are selected for the analysis of light curves because of their smallest incident angle to the direction of GRB 230307A. These two detectors both operate in two readout channels: high gain (HG) and low gain (LG), which are independent in terms of data processing, transmission, and dead-time. At the time of this burst, GECAM-C GRDs have a lower energy detection threshold of about 6 keV (owing to less radiation damage on SiPM) while GECAM-B GRDs have a relatively higher energy detection threshold of about 30 keV. For GRD04 of GECAM-B, the energy range of HG channel data are used from about 30 keV to 300 keV while the energy range of LG channel data are used from about 300 keV to 1000 keV. For GRD01 of GECAM-C, only HG channel data are used with the energy range from 6 keV to 30 keV. Though the response of GRD01 of GECAM-C for 6-15 keV and GRD04 of GECAM-B for 300-700 keV is affected by the electronics, this does not have any effect on the analysis of light curves. The background of GECAM-B is estimated by fitting the data from -50 s to -5 s and +160 s to +200 s with the first order polynomials. The background of GECAM-C is estimated by fitting the data from -20 s to -1 s and +170 s to +600 s with a combination of the first and second-order exponential polynomials[20]. The observed light curves are not directly used to identify features of broad pulse and fast pulses. Those features are identified through jointly fitting the background and the signals using various functional forms to describe pulses (Norris05, new FRED, etc.).
We have also cross-checked the GECAM data with that of Fermi/GBM, which has suffered from pile-up effects, which severely changed the shape of the light curve around the peak. For other part of the GBM data that did not suffer from pile-up, we performed the same analysis, including temporal analysis and spectral analysis. The results show well consistency with the GECAM results. We therefore do not include the GBM data in our following analysis.
1.2 Time-binned lightcurve and broad pulse fitting
We divide the data into eight energy bands (1st to 8th channels), namely, 6-30 keV, where the data are from GECAM-C; 30-70 keV, 70-100 keV, 100-150 keV, 150-200 keV, 200-300 keV, 300-500 keV and 500-1000 keV, where the data are from GECAM-B. The counts are binned every one second. The multiband count rates are then fitted to both the FRED pulse profile with Norris05 and our new FRED function (equations 1,2) superimposed on a background function. The fittings are done with the maximum likelihood method, which can be expressed as by assuming a Poisson distribution of the observed count rate, where the subscript i runs over all time bins from -4 s to 100 s (excluding 0 s to 1 s and 16 s to 20 s) and background part (as described in Observations part of Methods). The posteriors of the fitted parameters are found with a Monte Carlo Markov Chain (MCMC) method. The fitted parameters are listed in extended data Table 1.
In the Figure 2, we plot the the FRED parameters, namely and of Norris05 profile and of the new FRED profile, as functions of energy. The vertical error bars are the uncertainties of the fitted parameters, and the horizontal error bars indicate the width of each band. We fit power law relations between the FRED parameters and their corresponding energies. The power law fitting include the points from the 1st to the 6th channels (6-300 keV). The power law indices are -0.38, -0.37 and -0.36 for , and relations respectively. The uncertainties of the power index are found with Monte-Carlo samplings of the FRED parameters and from their distributions. The distributions of the FRED parameters are assumed to be Gaussian with their fitted uncertainties, while the distribution of energy is the corresponding energy deposition spectrum in each band.
| Norris05 | New FRED | ||||||
| Energy range | norm | t | norm | t | |||
| (keV) | (s) | (s) | (countss) | (s) | (s) | (countss) | (s) |
| 6-30 | |||||||
| 30-70 | |||||||
| 70-100 | |||||||
| 100-150 | |||||||
| 150-200 | |||||||
| 200-300 | |||||||
| 300-500 | |||||||
| 500-1000 | |||||||
1.3 Profile self-similarity checking
In the Extended Data Fig 1a, we stack the time re-scaled smoothed light curves, to show that the broad features of the light curves in different energy bands share the identical shapes. The smoothed light curves are convolution between a Gaussian kernel and the original light curves. The sigma of the Gaussian kernel is 1.5 s. We use the smoothed light curve in the 500-1000 keV band as the template, and we fit the template to light curves in other bands, which are scaled by multiplying scaling factors to the their time argument . It is intuitively shown that the self-similarity of the broad profile, and this conclusion is FRED formulation independent. The scaling factor as a function of energy is plotted in the Extended Data Fig 1b, which again shows a power law dependence.
[…]
b, The scaling factor of the Gaussian smoothed light curve as a function of energy. The vertical error bars indicate the 1 uncertainties of the fitted parameters, while the horizontal error bars indicate the ranges of the energy bins.
1.4 Fast pulses
After the FRED profiles are removed from the multi-band light curves, the residuals show fast pulses (Figure 1b). It is shown clearly that the fast temporal structures are aligned in time across all energy ranges. Here, we further test this temporal alignment with cross-correlation between the residuals. In the Figure 1c, we plot the normalized correlation coefficient as a function of lag time between the residuals in a certain band and that in band 6-30 keV. As we can see, all the cross-correlation peak at lag time zero, indicating the alignment of the fast pulses.
Here we further demonstrate that some of the prominent fast structures can be identified as small pulses. For instance, we show that the spike at s can be fitted with a pulse profile. We employ our new FRED formulation to fit the pulse from 3 s to 3.75 s, and find the typical width- relation ( in the new FRED formulation) in GRB pulses. In order to show this conclusion is independent on the pulse profile formulation, we use Gaussian function as an alternative profile to fit the spike, and obtain the same width- relation ( with the Gaussian pulse formulation. See the lower panel of Extended Data Fig 2b). It can be seen in the Extended Data Fig 2a that the FRED profile results in a better fit than the Gaussian profile. The fitted parameters are list in Extended data Table 2. It is also interesting to observe that this small pulse does not show spectrum lag of its peak time as the broad pulse. We demonstrate this by cross-correlating between the small pulses in multi-bands and that in 30-70 keV channel. The lag time together with their uncertainties as a function of energy is shown in the upper panel of Extended Data Fig 2b, where all lag times are almost zero, and we can clearly conclude that the lag time does not depend on energy like the broad pulse.
| Gaussian | New FRED | |||||
| Energy range | norm | norm | t | |||
| (keV) | (s) | (s) | (countss) | (s) | (countss) | (s) |
| 30-70 | ||||||
| 70-100 | ||||||
| 100-150 | ||||||
| 150-200 | ||||||
| 200-300 | ||||||
| 300-500 | ||||||
| 500-1000 | ||||||
[…]
The dip at s is another noticeable feature of the multi-band light curves. A natural attempt is to attribute such a dip to some sort of absorption or geometrical blocking. We define the effective “optical depth" in different energy bands as:
where is the net count rate at the bottom of the dip, which is found with a negative Gaussian fitting superposed on the FRED of broad pulse fitting from 17 s to 19.5 s (Extended Data Fig 3a); and is the net count rate of the broad feature. The results of the fitting with a negative Gaussian is tabulate in Extended data Table 3. It is intriguing to find that there is a power law energy dependence of the effective optical depth, and the best fit power law index is (Extended Data Fig 3c). This energy dependence of the optical depth challenges the absorption picture, for there is no known absorption mechanism whose cross section proportional to energy to the order of with a break at about 300 keV.
[…]
Here we view the dip as gaps between two successive fast temporal structures. Therefore, we fit a pulse profile (new FRED formulation) and a rising edge of another pulse (Norris05) to the dip at each energy bands from 17 s to 19.5 s (Extended Data Fig 3b). The fitted parameters are list in Extended data Table 4. The time scale of the earlier pulses as a function of energy is plotted in Extended Data Fig 3d, which shows clearly a typical power law energy dependence of pulse width.
| Energy range | Optical Depth | |||
| (keV) | (s) | (s) | (countss) | |
| 6-30 | ||||
| 30-70 | ||||
| 70-100 | ||||
| 100-150 | ||||
| 150-200 | ||||
| 200-300 | ||||
| 300-500 | ||||
| 500-1000 |
| pluse before gap | pluse after gap | additional constant | |||||
| Energy range | t | norm | t | norm | constant | ||
| (keV) | (s) | (s) | (countss) | (s) | (s) | (countss) | (countss) |
| 6-30 | |||||||
| 30-70 | |||||||
| 70-100 | |||||||
| 100-150 | |||||||
| 150-200 | |||||||
| 200-300 | |||||||
| 300-500 | |||||||
| 500-1000 | |||||||
1.5 Physical implications
Here we demonstrate that the observed features presented in this work cannot be explained within the framework of the internal shock (IS) model, but can be naturally reproduced with models involving many mini-jets in a single expanding magnetized fluid, as envisaged in the ICMART model.
In the standard IS picture [5], a random distribution of the shell bulk Lorentz factor results in internal shocks at random but similar distances from the central engine. These collisions generate a series of independent fast pulses. In this case, any long term overall trend of the light curve would have to reflect the activity history of the central engine. The same time history would be reflected in different energy bands. Even if each short pulse may have a energy-dependent behavior in the timescale of the short pulse itself, there is no reason to show a global behavior of shape-energy dependence (see the cartoon in Extended Data Fig. 4 for an illustration). To further demonstrate this, we conducted a series of simulations within the framework of the standard IS model with the following setting: We simulated the ejection of 50 shells from the central engine, with their initial thicknesses, Lorentz factors and masses sampled from random distributions in the logarithmic space. The Lorentz factors vary from 100 to 1000, the mass ranges from to , and the initial thickness ranges from to , all in cgs units. The central engine’s ejection times are randomly selected from a linear distribution, with values ranging from to seconds in the rest frame of the GRB central engine. Upon each collision, two shells merge, and the merged shell parameters (, , ) are adopted as the new values for the remaining shell. To ensure tracking future collisions, the “effective" ejection time of this new shell is redefined as , where and is the Lorentz factor of the merged shell. The simulation then restarts with one less shell. This process is repeated for each collision, allowing the code to monitor all collision/merging events for any arbitrarily designed central engine activity [6, 39]. The results of the simulation is presented in Extended Data Fig. 7a, which confirm our expectation that there is no shape-energy dependence of the overall lightcurve. In the IS model, the energy-dependent lag and width generally apply to each pulse. Even if there is a stretching factor in low energies to make pulses broader, the stretching factor scales with the duration of short pulses. There is no way to reproduce a large stretching factor in the scale of the broad pulse by superposing these short pulses.
In an variant IS model[7, 40, 41], a separate shock at a larger distance () is generated to make a slow pulse component that overlaps with those fast pulses. The overall lightcurve is dominated by the slow pulse and shows the shape-energy dependence (as illustrated in Extended Data Fig. 5). However, in this case, there should be an underlying slow pulse emission component, which is in direct contradiction with the observations presented in this work. The corresponding simulation results are presented in Extended Data Fig. 7b, which does not reproduce the observations, especially the deep dip at s. More details of the above-mentioned two IS simulations will be reported in a separated paper.
Within the framework of the ICMART model, collisions between a pair of highly-magnetized shells with different Lorentz factors would distort ordered magnetic field and trigger fast magnetic seeds, which will induce relativistic magnetohydrodynamic (MHD) turbulence. Numerical simulations have demonstrated that MHD turbulence has the ability to generate random local magnetic reconnection events [42, 43, 44]. These reconnection events would convert magnetic energy in the fluid to the particle energy, which would further distort magnetic fields, resulting in a reconnection cascade. The particle energy would then be transfer to the kinetic energy of the bulk through adiabatic cooling and to the observed photon energy through synchrotron radiation. The overall trend of the light curve corresponds to the growth of the number of magnetic reconnection events and a global curvature effect due to high-latitude emission after the line-of-sight emission ceases[15, 16]. Each mini-jet emission contributes to a fast pulse and their random latitudes and orientations cause the high variability of the lightcurve (see illustration in Extended Data Fig. 6). Dips could occur at random instances when the waiting time of pulses from successive mini-jets is long enough.
As the fluid expands, the gloabl magnetic field strength in the emission region naturally decreases. This leads to a global decline of synchrotron peak energy and causes a well-defined shape-energy dependence behavior, including spectral lags and global hard-to-soft evolution[17, 18]. Furthermore, photons from reconnection events at high latitudes arriving at later times after the line-of-sight emission is over lead to a redder spectrum because of its smaller bulk Doppler factor[16]. All these factors ensure an overall trend of a softer-wider/softer-later behavior.
We have presented a series of simulations to test this picture. We assume a pair of magnetized shells collide and merge as one fluid after a collision at cm. The mass of the fluid is assumed to be g, the initial Lorentz factor is the initial magnetization factor is (i.e. the comoving-frame magnetic energy is times of the rest-mass energy of the fluid). The number of reconnection events increases by a factor of during the emission period and the total number of the reconnection events within the cone is assumed to be . The relatively small number of reconnection events was chosen to ensure the occurrences of dips in most of the realizations. Each reconnection event has a size of cm, a random latitude within and a random oriention within . In Extended Data Fig. 7 c, we show one realization of the simulations that the following features consistent with the observations: 1) the lightcurve can be decomposed of many short pulses related to mini-jet emission; 2) the overall shape of the lightcurve resembles a single pulse, but can possess a very deep dip; 3) the overall single slow pulse has a significant shape evolution across energy bands. This simulation demonstrates that the ICMART model can well reproduce the observed behavior of GRB 230307A. Detailed simulations will be presented in a separate paper.
[…]
b, The results for the modified internal shock model that invokes superposition a large-radius internal shock and many small-radii internal shocks. The overall light curve is dominated by the emission from the large-radius internal shock, with no hope to produce dips in the energy-dependent lightcurves;
c, The results for the ICMART model. Both aligned short pulses and a global shape-energy dependence can be repduced. The shaded area indicates the location of a reproduced lightcurve dip due to the lack of mini-jet emission at that time.
Data Availability
The processed data are presented in the tables and figures of the paper, which are available upon reasonable request.
Code Availability
Upon reasonable requests, the code (mostly in Python) used to produce the results and figures will be provided.